OpenStudy (anonymous):

How the heck do I integrate the definite integral from 1 to 5 of s/(s^2+4)^1/2

OpenStudy (anonymous):

try \(u=s^2+4\)

OpenStudy (anonymous):

\[u=s^2+4, u(1)=5, u(5)=29, du = 2sds, \frac{1}{2}du=sds\] you get \[\int_5^{29}\frac{du}{\sqrt{u}}\]

OpenStudy (anonymous):

with at \(\frac{1}{2}\) out front that i forgot to write